OpenStudy (anonymous):

Let U = 5i+12j, V = -4i+j and W = i - 4j find: |2V-W|

OpenStudy (anonymous):

|2(-4i+j)-(i-4j)|

OpenStudy (anonymous):

yup! i got that

OpenStudy (anonymous):

but the answer came out to be root of 117 and I don't get how they got that :|

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

idk then, sorry

OpenStudy (anonymous):

it's fine thanks :D

OpenStudy (anonymous):

take the magniutde of the resulting vector

OpenStudy (anonymous):

I think i did the adding wrong or something would it be |-8i+2j-1-4j|

OpenStudy (anonymous):

and would that become |-9i+6j| ?

OpenStudy (anonymous):

yes... now just take the magnitude... sqrt((-9)^2 + (6)^2 )

OpenStudy (anonymous):

you can't root a negative right ?

OpenStudy (anonymous):

(-9)^2 = 81

OpenStudy (anonymous):

(-9)^2 = (-9)*(-9) = 81

OpenStudy (anonymous):

Ohhh i keep punching it in the calculator and it came out to be -81 but now that i think of it, negative x negative = positive, stupid me =.=

OpenStudy (anonymous):

thanks @dpalnc

OpenStudy (anonymous):

very common.. just so you know, the calculator does order of operations... if you entered -9^2, the calculator will do 9*9 and then take the opposite of the answer.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Ok!